The average (arithmetic mean) of the test scores of a class of "p" students is 70, and the average of the test scores of a class of "n" students is 92. When the scores of both classes are combined the aveage is 86. What is the value of p / n?
Answer by stanbon(48523)

The average (arithmetic mean) of the test scores of a class of "p" students is 70,
Total # of points for these p students is 70p
------------------------------------------------------
And the average of the test scores of a class of "n" students is 92.
Total # of points for these n students is 92n
------------------------------------------------------
When the scores of both classes are combined the aveage is 86.
Total # of points for the combined classes is 86(n+p)
------------------------------------------------------
What is the value of p / n?
----------------------------
Equation:
Grand total points = Grand total points
70p + 92n = 86(n+p)
70p + 92n = 86n + 86p
16p = 6n
p/n = 6/16
p/n = 3/8
===================
Cheers,
Stan H.